OpenStudy (anonymous):

Find ∫ 6√x dx, and hence evaluate

OpenStudy (anonymous):

\[\int\limits_{4}^{1}6\sqrt{x}dx\]

OpenStudy (anonymous):

My book says the answer is \[4x \sqrt{x}; 28\]

OpenStudy (mathmale):

Rewrite the definite integral replacing the sqrt(x) with \[x ^{\frac{ 1 }{ 2 }}.\]

OpenStudy (mathmale):

Your integral will now look like \[6 \int\limits_{}^{}x ^{\frac{ 1 }{ 2 }}dx\]

OpenStudy (mathmale):

Use the Power Rule for Integration.

OpenStudy (tkhunny):

Are you sure it's not [1,4],rather than [4,1]?

OpenStudy (anonymous):

I mean \[\int\limits_{1}^{4}6\sqrt{x}dx\]

OpenStudy (astrophysics):

\[\int\limits_{1}^{4} 6\sqrt{x} dx \implies \int\limits_{1}^{4} 6x^{1/2} dx \] Taking the integral you get \[\huge \implies \frac{ 6x^{1/2+1} }{ {1/2+1} } |_1^4\]

OpenStudy (astrophysics):

I leave the simplification to you :)